OpenStudy (anonymous):

taking the lim of n^12/ ( n^13- (n-3)^13) as n goes to infinity Cannot use derivatives . It is a sequence

OpenStudy (anonymous):

I cannot use derivatives or hopital's rule inthis problem

OpenStudy (anonymous):

@hartnn this is the question. I am trying the squeeze theorem to see if I can get something out of it. I know the limit is 1/39

OpenStudy (anonymous):

First answer to :(n-3)^13=?

OpenStudy (anonymous):

\[ lim_{ n\rightarrow \infty }\frac{n^{12}}{n^{13}-(n-3)^{13}}\]

OpenStudy (anonymous):

What is it ???

OpenStudy (anonymous):

@E.ali what do you mean what is it? It is a sequence

OpenStudy (anonymous):

Is it your question ?:)

OpenStudy (anonymous):

huh ???

OpenStudy (anonymous):

my question is how to prove the limit is 1/39 .

OpenStudy (watchmath):

use newton binomial expansion

OpenStudy (anonymous):

There's a difference between "showing" the limit is 1/39 and "proving" it's 1/39. The first would involve algebraically manipulating the expression so that finding the limit is simpler, and the second would involve showing that for some \(\epsilon>0\) there is some \(N\) such that \(n\ge N\) implies \(\left|\dfrac{n^{12}}{n^{13}-(n-3)^{13}}-\dfrac{1}{39}\right|<\epsilon\). Which one is it?

OpenStudy (anonymous):

@SithsAndGiggles i. YOu are right. The question says " evaluate the limit or prove it doesn't exists" So I guess showing it is 1/39 is enough

OpenStudy (anonymous):

@watchmath it worked I expanded and simplified the denominator , then factor n^12 and took the limit so the denominator went to 3*(13+0+0.......+0) =39 hence 1/39